### v7

Abstract
Lane detection is a critical component of Advanced Driver Assistance Systems (ADAS). Existing lane detection algorithms generally perform well under favorable weather conditions. However, their performance degrades significantly in adverse conditions, such as fog, which increases the risk of traffic accidents. This challenge is compounded by the lack of specialized datasets and methods designed for foggy environments. To address this, we introduce the FoggyLane dataset, captured in real-world foggy scenarios, and synthesize two additional datasets, FoggyCULane and FoggyTusimple, from existing popular lane detection datasets. Furthermore, we propose a robust Fog-Enhanced Network for lane detection, incorporating a Global Feature Fusion Module (GFFM) to capture global relationships in foggy images, a Kernel Feature Fusion Module (KFFM) to model the structural and positional relationships of lane instances, and a Low-level Edge Enhanced Module (LEEM) to address missing edge details in foggy conditions. Comprehensive experiments demonstrate that our method achieves state-of-the-art performance, with F1-scores of 95.04 on FoggyLane, 79.85 on FoggyCULane, and 96.95 on FoggyTusimple. Additionally, with TensorRT acceleration, the method reaches a processing speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its real-time capabilities and robustness in foggy environments.
keywords:
[…]
1 Introduction
Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents. As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving(Olovsson et al., 2022).
Under good weather and clear lane markings, current lane detection algorithms perform well. However, in complex road environments and poor weather, such as fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging. In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences. Ensuring safe driving in these conditions calls for the development of high-precision, highly robust lane detection methods optimized for foggy environments. As shown in Fig. 1, the in-vehicle camera captures real-time road images, which are preprocessed by edge devices and then sent to the lane detection module. Detected lane information is fused with data from other sensors, such as radar and GPS, along with real-time traffic data from the cloud. This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
Although end-to-end autonomous driving technologies have made significant progress at Level 3 (L3) and higher, most current vehicles, particularly commercial and urban vehicles, are still at Level 0 (L0), Level 1 (L1), and Level 2 (L2). This is partly due to the limited development of connected vehicle infrastructure and data platforms, which constrains the large-scale deployment of higher-level autonomous functions (Lim et al., 2021). In these lower levels of autonomous driving, lane detection remains a crucial technology, especially in complex and adverse road environments. Most existing vehicles rely on Advanced Driver Assistance Systems (ADAS) to enhance safety in complex scenarios, especially in active warning, lane-keeping, and other functions. Research related to vehicle trajectory analysis (Berghaus et al., 2024), congestion-aware traffic prediction (Zhang et al., 2025b), and 3D scene understanding (Qu et al., 2025) has become essential for improving system performance and safety in real-world driving environments, providing valuable insights for the further enhancement of lane detection capabilities in these challenging scenarios.
Figure 1: Lane detection in foggy scenarios for ADAS.(Pexels, 2025)
Earlier lane detection algorithms mainly relied on traditional computer vision methods, involving image preprocessing, feature extraction, and lane model fitting. Son et al.(Son et al., 2015) first detected the vanishing point to select the ROI adaptively, then extracted lane markings based on color features and performed clustering. Hou et al.(Hou et al., 2016) used the Canny operator for edge detection and applied a progressive probabilistic Hough transform to detect lines. While these feature-based methods are efficient and suitable for real-time applications, they require manual parameter tuning and struggle in challenging conditions such as worn or occluded lane markings.
With deep learning’s success in image tasks, it has become the mainstream approach for lane detection due to its strong feature extraction capabilities and better generalization. Deep learning offers higher robustness, adapting well to complex road environments and handling missing or occluded lane markings(Neven et al., 2018; Abualsaud et al., 2021; Yu et al., 2018; Pan et al., 2018). However, many networks rely on pixel-by-pixel classification, whereas lane detection does not require dense pixel segmentation. Qin et al.(Qin et al., 2020) addressed this by reframing lane detection as a row-based segmentation problem, reducing computational cost and improving detection speed.
In addition to segmentation methods, lane detection also includes detection-based approaches, many borrowing Anchor-based methods from object detection(Redmon, 2016; Farhadi and Redmon, 2018). Due to the elongated shape of lane lines, traditional bounding boxes are unsuitable. In lane detection, anchors are often preset starting points and lines with different slopes, with the network predicting the relative distance between the lane and the anchor to optimize shape and position(Chen et al., 2019; Zheng et al., 2022; Ran et al., 2023; Wang et al., 2025). These innovations improve accuracy and efficiency in complex or occluded scenarios.
Research on downstream tasks in adverse weather conditions, such as fog, is limited compared to other detection and segmentation tasks. Traditional solutions often use image enhancement techniques, like dehazing, to improve visual clarity, but this doesn’t always lead to better detection results. Some methods combine image enhancement with detection to reduce weather interference(Huang et al., 2020), while others simulate foggy images to augment datasets. For example, Nie et al.(Nie et al., 2022) generated foggy images from clear lane images in the CULane dataset, achieving synthetic data augmentation that improved recognition rates from 76.65% to 86.65% under light fog.
Domain adaptation techniques also address the style gap between foggy and clear-weather images(Chen et al., 2018; Zhang et al., 2021). Hnewa et al.(Hnewa and Radha, 2021) proposed a multi-scale domain-adaptive YOLO model, while Li et al.(Li et al., 2023) introduced a transfer learning framework for fog-adaptive object detection using labeled clear-weather and unlabeled foggy images, avoiding the need for annotated foggy datasets.
Traditional dehazing methods suffer from artifacts, whereas deep learning methods using CNNs capture more complex features but require high computational resources and diverse training datasets. However, obtaining real paired images under foggy conditions is challenging, and synthetic images often lead to reduced performance in real-world scenarios.
Lane detection is a critical component of Advanced Driver Assistance Systems (ADAS), supporting key functionalities such as lane departure warnings and autonomous navigation. While existing lane detection algorithms perform well under clear weather conditions with visible lane markings, their performance drops significantly in adverse environments like fog, where reduced visibility and image degradation obscure lane clarity, thus increasing the risk of traffic accidents. Furthermore, the lack of specialized datasets tailored to foggy conditions has limited the development of effective deep learning-based solutions for these challenging scenarios.
To address these limitations, we introduce the FoggyLane dataset, a new benchmark specifically designed for lane detection under foggy conditions. This dataset, comprising 1,423 annotated images across six distinct foggy road scenarios, fills a crucial gap in the current lane detection datasets. In addition, we generate foggy versions of two widely-used lane detection datasets, CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017), creating FoggyCULane and FoggyTusimple datasets. These datasets are essential for training and testing lane detection algorithms in foggy environments.
In response to the unique challenges posed by foggy conditions, we propose a robust real-time lane detection method optimized for such environments. The main contributions of our work are as follows:
  • 1)
    To address the issue of missing global information in foggy images, we design a Global Feature Fusion Module (GFFM) within the backbone network to capture relationships between inputs, improving feature extraction in foggy scenarios.
  • 2)
    To enhance the utilization of structural relationships between lane instances, we introduce a Kernel Feature Fusion Module (KFFM) that learns and predicts the correlations between lane instances, improving lane detection accuracy and robustness.
  • 3)
    To tackle the loss of edge information in foggy images, we incorporate a Low-level Edge Enhancement Module (LEEM) that enhances the model’s attention to edge features, improving the accuracy and stability of lane detection.
  • 4)
    To address the lack of open-source foggy lane detection datasets, we construct FoggyLane, a real-world dataset with 1,423 annotated images, and generate FoggyCULane and FoggyTusimple based on existing datasets. Our method achieves state-of-the-art performance with inference speeds of 192.9 FPS on NVIDIA RTX 3090 and 38.4 FPS on NVIDIA Jetson AGX Orin, confirming its real-time capability and practical applicability in real-world settings.
2 Related Work
In recent years, advances in deep learning have substantially improved lane detection performance. Building on these advances, current approaches for lane detection can be categorized into four main paradigms based on their modeling principles: segmentation-based, anchor-based, curve-based, and row-wise-based methods.
2.1 Segmentation-based Lane Detection
Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem. LaneNet(Neven et al., 2018) employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines. SCNN(Pan et al., 2018) extends the traditional layer-by-layer convolution operation with a slice-by-slice convolution operation, which passes the information horizontally and vertically between the pixels within a layer, thus enhancing the continuity of lane line recognition. RESA(Zheng et al., 2021) introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information. Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources. HW-Transformer(Zhao et al., 2024) presents an innovative lane detection network that utilizes row and column multi-head self-attention to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
2.2 Curve-based Lane Detection
Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations. PolyLaneNet(Tabelini et al., 2021b) introduces an end-to-end convolutional neural architecture that maps from images to polynomials representing each lane marker in the image, as well as domain lane polynomials and confidence scores for each lane. LSTR(Liu et al., 2021b) can also directly output the parameters of the lane shape model, has developed a Transformer-based network that takes into account non-local interactions, enabling it to capture the narrow structure of the lane lines and global context information. DBNet(Dai et al., 2024) introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation. However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection.
2.3 Anchor-based Lane Detection
Anchor-based approaches employ predetermined reference lines within the image space to generate lane predictions, subsequently calculating necessary offsets to precisely match these anchors with true lane markings. Line-CNN(Li et al., 2019) introduces the anchor mechanism into lane detection for the first time, and innovatively proposes line anchors. LaneATT(Tabelini et al., 2021a) effectively solves the problem of occlusion and illumination due to the lack of a line anchor, by fusing extracted anchor features with global features generated by attention module. CLRNet(Zheng et al., 2022) employs a coarse-to-fine detection strategy, where lane lines are first approximated using high-level semantic cues before being refined with detailed low-level features. This hierarchical fusion of semantic information significantly improves detection precision. FLAMNet(Ran et al., 2023) employs an adaptive line anchor strategy, augmented by patch-based feature pooling and decomposed self-attention mechanisms. This design strengthens both local feature extraction and global context modeling, improving performance in handling intricate lane structures. However, due to the multiple curvature and orientation variations of lane lines in the real world, the performance of the anchor-based methods may be limited in capturing precise paths with complex shapes.
2.4 Row-wise-based Lane Detection
Row-wise methods predict lane positions for each image row. E2E-LMD(Yoo et al., 2020) first proposes the row-wise-based lane detection task and designes an end-to-end architecture that directly outputs the lane marker positions. UFLD(Qin et al., 2020) utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes. CondLaneNet(Liu et al., 2021a) utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts. The row-wise-based detection methods improve computational efficiency by focusing only on the row regions in the image, and possesses high detection accuracy while maintaining efficiency.
3 Data Constrution
Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models. In the task of lane detection, models trained solely on clear weather data struggle to adapt to low-visibility foggy conditions. Issues such as blurred lane boundaries, faded lane colors, and reduced image quality often arise in fog, making lane detection more challenging. Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability. However, there is currently no publicly available foggy weather lane detection dataset. To address this gap, we have collected and constructed a foggy weather lane detection dataset called FoggyLane. In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017) datasets to test the generalizability of our methods.
Figure 2: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.(GoogleMaps, Accessed: 2024)
Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.(GoogleMaps, Accessed: 2024)
3.1 FoggyLane
Existing lane detection algorithms for intelligent connected vehicle ADAS systems perform well under favorable weather conditions, but their performance significantly deteriorates in adverse conditions like fog, which reduces visibility and increases the likelihood of traffic accidents. Current benchmark datasets, including CULane(Pan et al., 2018), Tusimple(Tusimple, 2017), and LLAMAS(Behrendt and Soussan, 2019), lack foggy weather data. To address this, we developed FoggyLane, a dataset specifically designed for foggy conditions. Data was collected using a monocular forward-facing camera at various times and locations, including Suzhou and Urumqi, China, between October 2023 and January 2024. The dataset includes 1,086 images at various resolutions, which were resized to $1640\times 590$ for consistency.
Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos(YouTube, Accessed: 2024), resulting in an additional 337 images. The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad. We split the dataset into training and testing sets at a 2:1 ratio, ensuring balanced distribution across scene types.
Comparing FoggyLane with existing datasets, such as Caltech-Lanes(Aly, 2008) and Tusimple(Tusimple, 2017), which focus on single traffic scenarios, and LLAMAS(Behrendt and Soussan, 2019), which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios. While ApolloScape (Huang et al., 2018) and BDD100k (Yu et al., 2020) include multiple traffic scenarios, they lack instance-level annotations. Similarly, although CULane (Pan et al., 2018) and CurveLanes (Xu et al., 2020) cover a wide range of traffic scenes, they do not include foggy conditions. Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced. To address this gap, FoggyLane provides a comprehensive solution for foggy weather lane detection with instance-level annotations, available in both Tusimple (Tusimple, 2017) and CULane (Pan et al., 2018) formats for broader accessibility.
Figure 4: The different road types of FoggyLane images examples.
Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes (Aly, 2008) 1.2k None 640 x 480 Tusimple (Tusimple, 2017) 6.4k None 1280 x 720 LLAMAS (Behrendt and Soussan, 2019) 100k None 1276 x 717 Appolloscape (Huang et al., 2018) 110k None 3384 x 2710 BDD100k (Yu et al., 2020) 100k None 1280 x 720 CULane (Pan et al., 2018) 133k None 1640 x 590 CurveLanes (Xu et al., 2020) 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
Our dataset has two key advantages:
  1. []
  2. 1.
    Pioneering foggy lane detection dataset: FoggyLane is a dataset specifically designed to address foggy weather scenarios. It covers a range of challenging conditions, including foggy congested roads, foggy curves, and foggy night roads. This dataset is crucial for testing the robustness of ADAS under adverse conditions.
  3. 2.
    Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.(Zhang et al., 2024a). These rigorous annotations can significantly improve the accuracy and consistency of lane detection in foggy weather, thus enhancing the stability and safety of autonomous driving systems and providing a solid data foundation for the reliable deployment of autonomous technology.
Figure 5: Networks of foggy image generation. $\otimes$ represents element-wise multiplication.
Figure 6: The FoggyCULane dataset examples.
3.2 FoggyCULane and FoggyTusimple
Currently, numerous datasets are available for lane detection. Among them, we select two widely adopted and representative benchmarks: CULane(Pan et al., 2018) and TuSimple(Tusimple, 2017). The CULane dataset (Pan et al., 2018) is a large-scale and challenging benchmark that covers diverse scenarios including highways, urban streets, and rural areas in Beijing. In contrast, the TuSimple dataset (Tusimple, 2017) primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations. Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
The method for generating foggy images is inspired by the model (Nie et al., 2022), and the complete generation process is shown in Fig. 5. To be specific, the atmospheric scattering model can effectively utilize prior information to analyze the degradation mechanisms of image quality in foggy conditions. It models foggy images as a combination of scattered light and direct light based on optical principles, ultimately synthesizing a clear image into a foggy one. The mathematical modeling equation defined by Koschmieder (Harald, 1924) is represented as:
$I(x)=J(x)t(x)+A(1-t(x))$ (1)
where $x$ indicates a certain pixel of the image, $J(x)$ refers to the original clear image of the object and $I(x)$ refers to the foggy image. $A$ is the atmospheric light value at infinity and $t(x)$ can be further defined as:
$t(x)=e^{-\beta d(x)}$ (2)
where $\beta$ is the extinction coefficient, which is closely related to the concentration, size, type and distribution of particulate matter. $d(x)$ represents the distance between the object and the viewer. As this distance increases, the transmittance decreases, resulting in thicker fog.
Figure 7: The FoggyTuSimple dataset examples. (a) is the original TuSimple dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
To calculate the corresponding $d(x)$, we use the S2R-DepthNet(Chen et al., 2021) that has good generalization performance across various applications. The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image. Then they suppress information irrelevant to depth and use structural information for depth prediction. The Structure Extraction Module (STE) decomposes the image to obtain domain-invariant structure information. The Depth-specific Attention (DSA) predicts an attention map that helps to suppress depth-irrelevant structures by leveraging high-level semantic information extracted from the raw image. After passing through the DSA Module and STE Module, as shown in Fig. 5, the attention map and the structure map are multiplied to obtain a concise and comprehensive depth-specific structure map. The Depth Prediction (DP) is responsible for depth estimation.
According to the transmittance formula Eq. (2), we use the dark channel prior(He et al., 2011) to calculate the average atmospheric light value $A$. In the non-sky area of an outdoor clear image, there is usually at least one very small or close to zero value in the RGB image for each pixel, which is called the dark channel of that pixel. Therefore, for any image $I$, the mathematical formula for its corresponding dark channel image $I_{\text{dark}}(x)$ can be defined as:
$I_{\text{dark }}(x)=\min_{c\in R,G,B}\left(\min_{y\in\Omega(x)}I_{c}(y)\right)$ (3)
where $I_{\text{c}}(y)$ is the value of the c-th channel of image $I$ at pixel point $y$, and $\omega(x)$ is a small region around pixel point $x$. In the dark channel, find the top 0.1% brightest pixels and take the maximum value of their RGB channels as an approximation of the average atmospheric light value $A$. After obtaining the transmittance and the average atmospheric light value, the images of light fog, medium fog, and heavy fog can be obtained by fine-tuning the extinction coefficient $\beta$. The original images and the foggy images generated under different $\beta$ values are displayed in Figs. 6 and 7.
4 Methods
4.1 Overall structure
To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose a fog-enhanced lane detection network inspired by row-wise methods. As illustrated in the Fig. 8, the network comprises three main components: a backbone network, a neck network, and a detection head. To mitigate the loss of global information in foggy images, we introduce a global feature fusion module within the backbone network. In the neck network, we utilize HRNetV2(Wang et al., 2021) to further integrate multi-scale features from different stages of the backbone, enabling the aggregation of information across various scales and resolutions.
For the detection head, we employ the concept of dynamic convolution, incorporating both lane instance detection and lane mask segmentation. During instance detection, the network identifies lane instances within the image and predicts a set of dynamic convolution kernel parameters for each lane. These parameters are subsequently applied to the lane mask segmentation head to filter the instance-specific segmentation map for each lane. This approach eliminates the need for complex post-processing steps, enhancing detection speed.
Moreover, to fully exploit the structural relationships between lane lines, we designed a kernel feature fusion module, which allows the network to automatically learn the correlations between predicted lane instances. This contributes to more accurate lane identification and localization. Additionally, to compensate for the loss of edge information in foggy images, we incorporated a low-level edge enhancement module, further improving the network’s ability to detect lane boundaries with precision.
Figure 8: Overview of the proposed network.
Figure 9: Backbone Network with Global Feature Fusion.
4.2 Global Feature Fusion Module
In recent years, Transformer models have achieved significant success in natural language processing, leading researchers to explore their applications in computer vision. With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information. As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks(Zhang et al., 2025a; Chen et al., 2023; Zhang et al., 2024b). However, applying Transformers to the visual domain still faces some challenges. Firstly, traditional Transformer architectures lack the hierarchical structure and down-sampling operations found in CNN models, making them less adaptable to changes in resolution and receptive field, which may hinder handling of complex visual entities. Additionally, calculating global attention at high resolutions can be computationally intensive.
To address these issues, researchers have proposed several enhanced architectures, such as hierarchical Transformers(Liu et al., 2021c, 2022; Dong et al., 2022). Among them, the Swin Transformer(Liu et al., 2021c) combines CNN-inspired concepts, integrating local perception with global attention in a hierarchical Transformer framework. As the Swin Transformer has shown strong performance in classification, detection, and segmentation tasks, balancing computational efficiency with accuracy, we chose the Swin Transformer as our backbone network. However, since the Swin Transformer’s window attention mechanism primarily focuses on local features, its hierarchical structure may still lack sufficient global context for foggy scene understanding. Therefore, as illustrated in the Fig. 9, we designed a Global Feature Fusion Module (GFFM) to enhance the network’s global information integration. Specifically, we replaced the window attention mechanism in the final stage with a standard multi-head attention mechanism to capture global correlations across all inputs. At this stage, image resolution has been sufficiently reduced, making the computational cost of global attention manageable.
4.3 Kernel Feature Fusion Module
Dynamic convolution adjusts convolutional kernel parameters based on input data, enhancing network flexibility and generalization compared to traditional convolution. CondInst(Tian et al., 2023) is an instance segmentation framework using dynamic convolution to generate kernel parameters for each instance, reducing network load and improving efficiency. Similarly, CondLane(Liu et al., 2021a) applies dynamic convolution for lane instance segmentation, achieving strong performance.
Building on this, our lane detection head also uses dynamic convolution, consisting of lane instance detection and mask segmentation. Unlike traditional methods, we generate dynamic convolution kernels using the lane’s starting position features rather than center features, which are more suitable for the elongated nature of lanes. We design an instance detection head with two branches: one predicts lane starting points and generates a heatmap, and the other generates dynamic convolution kernels for lane segmentation.
However, existing methods don’t fully utilize the structural relationships between lane instances. To address this, we introduce a Kernel Feature Fusion Module (KFFM) that enables the network to learn and predict relationships between lane instances, enhancing detection performance. KFFM is shown in the bottom right of Fig. 8.
4.4 Low-level Edge Enhanced Module
Images in foggy conditions differ from those captured under good lighting primarily due to their blurriness and low contrast. This effect renders the edges of objects and details within the scene indistinct and hard to detect, impacting the algorithm’s ability to understand and recognize the scene accurately. Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking. Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig. 10. It is evident that fog significantly degrades edge information, particularly affecting lane line edges, which become nearly indiscernible. Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—both global scene information and detailed edge information are essential for accurate lane detection.
Figure 10: Comparison of edge information between foggy and normal scenarios.
To address the challenges of lane detection in foggy conditions, we apply global feature fusion within both the backbone network and detection head to enhance high-dimensional global features. Additionally, we use multi-scale fusion to aggregate detail features from low-resolution feature maps. However, this approach alone does not fully compensate for the loss of edge detail inherent in foggy images. Therefore, we introduce a Low-level Edge Enhanced Module (LEEM) into the network to boost the network’s focus on edge information, thereby improving both the accuracy and stability of lane detection.
Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN(Pan et al., 2018) AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane(Liu et al., 2021a) ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet(Zheng et al., 2022) CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet(Ran et al., 2023) IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 91.99 84.23 68.08 35.27 34.38 18.60 70.0 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 91.07 79.38 57.44 14.74 52.42 19.04 171.6 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 90.02 83.44 71.66 40.51 51.97 16.05 150.4 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
  • Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine. * indicates the FPS after acceleration with TensorRT.
Specifically, we add an auxiliary path in the network to learn the edge features of the input image. First, we apply the Canny operator to the input image for edge detection, supplemented by lane line labels to reconstruct any missing lane edges, generating the training labels for the edge enhancement module. To avoid interfering with high-level semantic extraction and subsequent detection segmentation, while strengthening the low-level network’s learning of edge textures and other low-dimensional features, we introduce a branch at the second stage of the backbone network. This branch sequentially applies a 3 × 3 and a 1 × 1 convolution. We then compute the cross-entropy loss between the branch output and the edge labels. The detailed structure is illustrated in the bottom-left of the Fig. 8. Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
Figure 11: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyLane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
4.5 Loss functions
Based on the above, our method involves a total of five prediction tasks: lane start point prediction, row-based segmentation prediction, vertical range prediction, offset prediction, and edge prediction. The loss functions for each task are detailed below.
4.5.1 Lane start point prediction
In feature maps with a large receptive field and low resolution, the lane line start point is difficult to be accurately represented by a single pixel location. Moreover, in such large receptive field feature maps, pixels near the target point exhibit similar features. If these pixels are directly labeled as negative samples, it will interfere with the network training and make it difficult for the network to learn and optimize. Therefore, this method uses a Gaussian kernel function to generate a heatmap label for the lane line start point. Each element in the heatmap label represents the confidence that the position corresponds to the target point; the closer to the target point, the higher the confidence, and the farther away, the lower the confidence. In other words, the ground truth is softly labeled with a Gaussian function, making it easier for the network to converge. Suppose $(x,y)$ denotes a position on the low-resolution feature map, and $(p_{x},p_{y})$ represents the coordinate of the target point on the low-resolution feature map. Then, the ground truth heatmap $GT$ value at position $(x,y)$ is defined as:
$GT(x,y)=\exp\left(-\frac{(x-p_{x})^{2}+(y-p_{y})^{2}}{2\sigma^{2}}\right)$ (4)
where $\sigma$ controls the decay rate of the Gaussian kernel. A larger $\sigma$ results in slower confidence decay and thus a larger heatmap area.
This method combines the Gaussian heatmaps of all lane line start points within the image into a single heatmap. When multiple heatmaps respond at the same location, the maximum value is retained to form the final ground truth heatmap. Then, the Focal Loss is used to calculate the loss between the predicted heatmap and the ground truth heatmap. Let the predicted heatmap value at position $(x,y)$ be $P_{xy}$, and let $N$ be the total number of pixels in the heatmap. The Focal Loss between the predicted heatmap and the ground truth heatmap is defined as:
$\mathcal{L}_{\text{hm}}=-\frac{1}{N}\sum_{x,y}\begin{cases}(1-P_{xy})^{\alpha}\log(P_{xy}),&\text{if }GT(x,y)=1\\ (1-GT(x,y))^{\beta}(P_{xy})^{\alpha}\log(1-P_{xy}),&\text{otherwise}\end{cases}$ (5)
4.5.2 Row-based segmentation prediction
Row-based segmentation predicts a horizontal coordinate $\mathrm{E}\left(\hat{x}_{i}\right)$ for each row. For this task, our method uses the L1 loss function to optimize the predicted coordinates. Suppose $V$ denotes the effective vertical range of the lane line, $N_{v}$ is the number of effective rows, and $x_{i}$ is the ground truth horizontal coordinate. The loss function is defined as:
$L_{\text{row}}=\frac{1}{N_{v}}\sum_{i\in V}\left|\mathrm{E}\left(\hat{x}_{i}\right)-x_{i}\right|$ (6)
4.5.3 Vertical range prediction
For the label of the lane line vertical range vector prediction, our method uses a $1\times Y$ vector representation. If the lane line passes through the $Y_{i}$ row, its value is 1; otherwise, it is 0. Thus, this can be regarded as a segmentation task. Our method optimizes it using the cross-entropy loss function. Suppose $v_{i}$ denotes the predicted probability that the $i$-th row is positive, then the loss function is defined as:
$L_{\text{range}}=\sum\left(-y_{gt}^{i}\log(v_{i})-(1-y_{gt}^{i})\log(1-v_{i})\right)$ (7)
4.5.4 Offset prediction
Firstly, the generation of ground truth labels for offset prediction is required. For each lane line, the proposed method constructs the ground truth offset as the horizontal displacement relative to the lane centerline for each grid cell within a specified width. The predicted offsets are optimized using an L1 loss function. Let $\delta_{xy}$ denote the ground truth offset at coordinate $(x,y)$, $\hat{\delta}_{xy}$ the predicted offset, and $\Omega$ the region surrounding the lane line. The loss function is then formulated as follows:
$L_{\text{offset}}=\frac{1}{N_{\Omega}}\sum_{(x,y)\in\Omega}\left|\hat{\delta}_{xy}-\delta_{xy}\right|$ (8)
4.5.5 Edge prediction
For the task of edge enhancement, the method treats it as an edge segmentation problem, which is essentially a binary classification task. The labels at edge locations are assigned a value of 1, while non-edge locations are assigned 0. Accordingly, the cross-entropy loss function is employed for optimization. Let $e_{gt}^{i}$ denote the label of pixel $i$, and $p_{i}$ represent the predicted probability of the pixel being positive. The loss function is defined as:
$L_{\text{edge}}=\sum_{i}\left(-e_{gt}^{i}\log(p_{i})-(1-e_{gt}^{i})\log(1-p_{i})\right)$ (9)
4.5.6 Total loss
To balance the influence of different tasks on the overall training of the network, a weighted fusion of the loss functions for each sub-task is employed. The total loss function is formulated as follows:
$L_{\text{total}}=\alpha L_{hm}+\beta L_{\text{row}}+\gamma L_{\text{range}}+\delta L_{\text{offset}}+\varepsilon L_{\text{edge}}$ (10)
where the network achieves optimal performance when the weights are set as $\alpha=1$, $\beta=1$, $\gamma=1$, $\delta=0.4$, and $\varepsilon=0.1$.
Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN(Pan et al., 2018) AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane(Liu et al., 2021a) ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet(Zheng et al., 2022) CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet(Ran et al., 2023) IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 90.85 73.32 67.18 75.32 47.18 87.14 63.38 1456 69.20 74.61 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 93.34 77.76 75.26 79.76 53.25 89.62 75.94 1289 74.54 78.68 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 92.90 77.50 74.03 80.89 51.71 89.76 73.02 1226 74.21 79.13 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.72$\uparrow$)
  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow, Curve and Night conditions, as well as in the overall (Total) performance.
Figure 12: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyCULane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
5 Experiments
5.1 Datasets
In the Data Construction chapter, we introduce the FoggyLane dataset, which was constructed from real-world collections for foggy lane detection, along with the FoggyCULane and FoggyTusimple datasets that were synthesized through modeling foggy images. Our method is evaluated on these three datasets.
FoggyLane is a lane detection dataset specifically designed for foggy scenarios, containing 1,423 frames of foggy lane images for training and testing. All images are of resolution 1640x590 pixels.
FoggyCULane is synthesized from the large-scale challenging lane detection dataset CULane using a foggy image modeling approach. It maintains the same number of images as CULane, which includes 88,880 images for training and 34,680 images for testing. The resolution of all images is also 1640x590 pixels.
FoggyTusimple is synthesized from one of the most widely used benchmark datasets for lane detection, Tusimple, through the foggy image modeling method. At the same time, the dataset follows the same structure as Tusimple, comprising 3,268 training images, 358 validation samples, and 2,782 test cases. All visual data maintains a consistent resolution of 1280×720 pixels.
Figure 13: Several metrics of our model’s performance during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
5.2 Evaluation Metric
Lane detection is commonly treated as a semantic segmentation task, evaluated using standard metrics such as True Positive (TP), False Positive (FP), False Negative (FN), and True Negative (TN). However, due to its unique structure, lane detection also adopts specialized metrics. Following SCNN’s(Pan et al., 2018) protocol, we connect labeled or predicted lane points into line segments and expand them into 30-pixel-wide areas. A prediction is considered correct if the Intersection over Union (IoU) with the ground truth exceeds 0.5. Based on this, we compute Precision, Recall, and F1-score, with the following formulas:
$F1=\frac{2\times Precision\times Recall}{Precision+Recall}$ (11)
$Precision=\frac{TP}{TP+FP}$ (12)
$Recall=\frac{TP}{TP+FN}$ (13)
For the FoggyTusimple dataset, we adopt the evaluation metrics of the Tusimple dataset. The official Tusimple dataset provides three standard evaluation metrics: false positive rate (FPR), false negative rate (FNR), and accuracy. The formula for calculating Accuracy is as follows:
$\textit{Accuracy}=\frac{\sum_{\textit{clip}}C_{\textit{clip}}}{\sum_{\textit{clip}}S_{\textit{clip}}}$ (14)
where $C_{\textit{clip}}$ represents the number of correctly predicted lane points in the image, while $S_{\textit{clip}}$ denotes the total number of ground truth lane points. For the pixels in the image, a predicted lane point is considered correct if it falls within 20 pixels to the left or right of the ground truth. Furthermore, if more than 85% of the predicted lane points on a single lane line correspond to the ground truth lane points, the predicted lane is regarded as a true positive.
Table 4: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN(Pan et al., 2018) AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet122 93.24 95.43 9.61 3.50 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CondLane(Liu et al., 2021a) ICCV 2021 ResNet101 96.03 95.14 2.94 5.09 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet101 96.13 95.64 4.15 3.58 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 95.54 95.91 4.32 4.60 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 96.33 95.91 4.50 2.77 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 96.37 94.99 3.21 4.09 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.62$\downarrow$) 3.83
5.3 Implementation Details
We conducted experiments on the NVIDIA GeForce RTX 3090 platform. The SwinGFFM backbone network was pre-trained on the COCO dataset for 500 epochs. During lane detection training, we loaded the pre-trained weights and resized input images to 800 × 320. Various data augmentation techniques, including rotation, flipping, brightness adjustment, and random scaling, were applied. The batch size was set to 16, with an initial learning rate of $3\times 10^{-4}$, using the Adam optimizer and a Step learning rate decay. For full network parameters and training details, please refer to our open-source code.
To benchmark against existing methods(Pan et al., 2018; Zheng et al., 2021; Tabelini et al., 2021a; Zheng et al., 2022; Ran et al., 2023; Liu et al., 2021b; Dai et al., 2024; Qin et al., 2020; Liu et al., 2021a; Wang et al., 2025), we followed the training settings provided in the respective papers. For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results.
5.4 Quantitative Results
5.4.1 Results on FoggyLane
We evaluated our proposed method on the FoggyLane dataset and compared it with other leading lane detection methods. As shown in the Tabel 2, we mainly compared the methods in terms of F1 scores, computational complexity (GFlops), processing speed (FPS).
From the Table 2, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively. These results surpass the best-performing alternative methods by margins of 2.95%, 3.42%, 3.94%, and 3.41%. The mean F1 (mF1) score demonstrates a 3.13% improvement over the strongest competing approach. Notably, the SwinGFFM-t variant achieves an F1@50 score of 92.34, which even exceeds the performance of DLA34-based CLRNet, previously the top-performing method among comparative approaches.
5.4.2 Results on FoggyCULane
We compared our method with other advanced lane detection methods on the FoggyCULane dataset. As shown in the Table 3, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve. Additionally, our model achieved the highest overall F1 score of 79.85,which achieves a 0.72% improvement over the best alternative method.
5.4.3 Results on FoggyTusimple
As shown in the Table 4, the traffic scenes in the FoggyTusimple dataset are relatively simple, resulting in smaller performance differences between methods. Nonetheless, our method achieved an F1 score of 96.95, outperforming the best existing method by 0.43%. This further demonstrates the effectiveness of our method in foggy conditions.
5.5 Ablation Study
We conducted ablation studies on the FoggyLane and FoggyCULane datasets to evaluate the effectiveness of the GFFM, KFFM, and LEEM modules, using Swin-t as the backbone. As shown in Tables 5 and 6, each module improves the model’s performance. The GFFM module contributes the most, increasing the F1-score by 1.07 on FoggyLane and 0.41 on FoggyCULane, highlighting its ability to capture both local and global information. Adding the KFFM module further boosts the F1-score by 0.63 on FoggyLane and 0.26 on FoggyCULane. Finally, the LEEM module improves edge detection, increasing the F1-score by 0.28 on FoggyLane and 0.14 on FoggyCULane.
Table 5: Ablation experiments of different components on the FoggyLane dataset.
  Model GFFM KFFM LEEM F1-score Gflops(G)
Baseline 90.36 13.77
91.43 11.94
92.06 11.97
Ours 92.34$\uparrow$ 12.01
Table 6: Ablation experiments of different components on the FoggyCULane dataset.
  Model GFFM KFFM LEEM F1-score Gflops(G)
Baseline 77.85 13.77
78.26 11.94
78.52 11.97
Ours 78.66$\uparrow$ 12.01
5.6 Model deployment
Figure 14: A edge computing system built with NVIDIA Jetson AGX Orin. After experimental validation, it has been demonstrated that our method can run accurately on the vehicle’s edge computing platform with limited computational resources, achieving a performance of 38.4 FPS.
To further evaluate the performance of our algorithm in real-world scenarios, in addition to experiments on existing datasets, we also deployed and tested the model on the vehicle. Specifically, we implemented an edge computing system on the vehicle, as illustrated in the Fig. 14. This system comprises the NVIDIA Jetson AGX Orin edge computing device, a monitor, a camera, and a power supply unit. The results of the evaluation demonstrate that our approach can operate in real-time with high accuracy on the vehicle’s edge computing platform, despite the constraints on computational resources.
6 Conclusion
This study focuses on lane detection in foggy conditions, aiming to address the range of challenges posed by foggy environments, which is crucial for ensuring driving safety in fog and enhancing Advanced Driver Assistance System (ADAS) capabilities. To address the current lack of publicly available datasets for lane detection in foggy conditions, we constructed a real-world foggy lane dataset, FoggyLane, and synthesized two additional foggy datasets, FoggyCULane and FoggyTusimple, based on the popular CULane and Tusimple datasets.
In response to the challenges of lane detection in foggy scenarios, we propose an end-to-end fog-enhanced lane detection method. First, this method incorporates a Global Feature Fusion Module in the backbone network to address the absence of global spatial information in foggy scenes. Second, we employ a dynamic convolution-based detection head for instance-sensitive lane segmentation, integrating a Kernel Feature Fusion Module within the detection head to automatically capture associations between lane instances. Additionally, a Low-Level Edge Enhancement Module was added to increase the network’s focus on edge information, mitigating the loss of edge details and improving lane detection accuracy under foggy conditions.
Comparative experiments on the FoggyLane, FoggyCULane, and FoggyTusimple datasets with other advanced lane detection methods demonstrate that our approach achieves superior performance in both clear and foggy conditions, significantly improving lane detection accuracy. Furthermore, with TensorRT acceleration, our method achieves an inference speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, further validating its practical applicability in real-world scenarios.
Acknowledgments
This project is jointly supported by National Natural Science Foundation of China (Nos. 52172350, W2421069 and 51775565), the Guangdong Basic and Applied Research Foundation (No. 2022B1515120072), the Guangzhou Science and Technology Plan Project (No. 2024B01W0079), the Nansha Key RD Program (No. 2022ZD014), the Science and Technology Planning Project of Guangdong Province (No. 2023B1212060029).We would like to release our source code and dataset.